# Stabilizing Rate of Stochastic Control Systems with Multiplicative Noise

### v1

@@ 6.2.1 Single Near-Marginal Instance
    The noise-free, uncontrolled part of the system already lies at the boundary of stability, since $A$ has eigenvalues very close to one, and even weak multiplicative noise can readily push the overall dynamics into an unstable regime.
@@ 6.2.2 Scaling Study Across the Stability Boundary
Objective and Methodology.
To stress-test RNVI beyond a single operating point, we smoothly vary the system across the marginal regime. We introduce a global scaling factor
@@
Results.
Table 2 summarizes the bounds for $\rho^{*}$ at $20$ values of $\theta$. The corresponding curves and bound width are visualized in Figures 56.
@@ Results.
Discussion.
Three points emerge clearly from the scaling study.
@@ Discussion.
Figure 5: Certified bounds for $\rho^{*}$ versus the scaling $\theta$.
@@ 7 Conclusion
We develop a quantitative framework for analyzing mean-square exponential stabilization of stochastic systems with multiplicative noise under state-feedback control policies. The study introduces the optimal stabilizing rate $\rho^{*}$ as a measure of how fast stabilization can be achieved, and pursues two complementary directions. On one hand, norm-based conditions are extended to the stochastic setting, providing a necessary and sufficient condition for the exact attainability of the optimal stabilizing rate $\rho^{*}$ and conservative upper and lower estimates. On the other hand, we reformulate the Problem (OC) into an optimal control Problem (OSR) and develope the RNVI algorithm, obtaining the computable two-sided bounds for the optimal stabilizing rate under any admissible state-feedback control policies, while also producing a near-optimal feedback gain.

### v2

@@ 6.2.1 Single Near-Marginal Instance
    The noise–free, uncontrolled part of the system already lies at the boundary of stability, since $A$ has eigenvalues very close to one, and even weak multiplicative noise can readily push the overall dynamics into an unstable regime.
@@ 6.2.2 Scaling Study Across the Stability Boundary
Objective and Methodology. To stress-test RNVI beyond a single operating point, we smoothly vary the system across the marginal regime. We introduce a global scaling factor
@@
Results. Table 2 summarizes the bounds for $\rho^{*}$ at $20$ values of $\theta$. The corresponding curves and bound width are visualized in Figures 910.
@@
[…]
Figure 8: Certified squared-gap $\Delta_{\tau}=\frac{\tau}{n(1-\tau)}\lambda_{\max}\big((P^{(\tau)})^{-1}\big)$ for each $\theta$.
@@
Figure 9: Certified bounds for $\rho^{*}$ versus the scaling $\theta$.
Discussion. Three points emerge clearly from the scaling study.
@@
  4. 4.
    Resolution near criticality and decay of the certified gap. In the near-critical parameter range, where both bounds for $\rho^{\ast}$ are close to one, the certified squared-gap $\Delta_{\tau}$ produced by RNVI decreases steadily as $\tau$ is reduced. This behavior is consistent with the relaxed tightness requirement described in Remark Remark: for the values of $\tau$ used in this example, the quantity $\tau\,\lambda_{\max}\!\big((P^{(\tau)})^{-1}\big)$ remains numerically small, and the resulting certificates sharpen accordingly. Thus, even when the stabilizing rate is very close to the transition between stability and instability, RNVI continues to provide reliable and increasingly tight two-sided bounds as $\tau$ is refined.
@@ 7 Conclusion
We develop a quantitative framework for analyzing mean-square exponential stabilization of stochastic systems with multiplicative noise under state-feedback control policies. The study introduces the optimal stabilizing rate $\rho^{*}$ as a measure of how fast stabilization can be achieved, and pursues two complementary directions. On one hand, norm-based conditions are extended to the stochastic setting, providing a necessary and sufficient condition for the exact attainability of the optimal stabilizing rate $\rho^{*}$ and conservative upper and lower estimates. On the other hand, we reformulate the Problem (OC) into an optimal control Problem (OSR) and developed the RNVI algorithm, obtaining the computable two-sided bounds for the optimal stabilizing rate under any admissible state-feedback control policies.